OpenStudy (anonymous):

What is the x-intercept and the y-intercept of the graph of 6x - 3y = 18? Show your work

OpenStudy (anonymous):

What do you know about the x and y intercepts?

OpenStudy (anonymous):

not so much lol

OpenStudy (anonymous):

Okay so the x intercept is where the line crosses the x axis and at this point, we know y must be equal to 0. Likewise, for the y intercept we know x=0. To calculate the y intercept, simply sub x=0 into the original equation and solve for y. For the x intercept, sub y=0 into the equation and solve for x. See if you can do this and let me know if you get stuck.

OpenStudy (superhelp101):

just like tom said your x intercepts are where y is 0 and y intercepts are where x is 0 so plug in like this \[6(0)-3y=18\] \[0-3y=18\] \[-3y=18 \] \[y=-6\] there fore your intercepts is (0,-6) Then for your x intercepts plug 0 in for y \[6x-3(0)=18\]\[6x=18\]\[x=3 \] therefore your x intercepts is (3,0)

OpenStudy (superhelp101):

The y intercepts is (0,-6) and x intercepts (3,0)

OpenStudy (superhelp101):

hope that makes sense :/ :D

OpenStudy (superhelp101):

\(\bbox[5pt, #33fd33 ,border:5px solid #872af0]{\bf\huge\I\ hope\ that helps!!}\)